# A defect-chemistry-informed phase-field model of grain growth in oxide electroceramics

### v2

@@ 6 Conclusion
    The proposed phase-field model reproduces a series of benchmark results at equilibrium of Fe-doped STO bicrystal. Simulation results, including the symmetric SCLs, grain boundary potentials, concentrations of oxygen vacancies and acceptor dopants in the GB core and the bulk phase as well as the electrostatic distributions, show good agreement with the analytical predictions in the MS and GC models.
@@
    Simulations are carried out comprehensively for a GB core moving with constant velocities under quasi-equilibrium state. At low core velocity, the accumulation zone in the shrinking grain decreases, while it increases in the growing grain, leading to the asymmetric SCL formation. With increasing GB core velocity, the dopant segregation diminishes. However, the movement of oxygen vacancy within the GB core only breaks at extremely high velocity due to the high diffusivity of oxygen vacancy. In addition, the electrostatic potential in the shrinking grain becomes higher than that in the growing grain.
@@
    Abnormal grain growth is observed even without distinguishing any grain orientation or considering any anisotropic GB mobility in the phase-field simulations at 1623K. It indicates that the solute drag effect alone can lead to bimodal grain size distribution. As the formation energy differences of oxygen vacancies and acceptor dopants ($\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$) become more negative, the grain growth velocity decreases. This trend is also observed with increasing dopant concentration ($c_{\text{dop},\text{b}}$). Notably, the dopant concentration profile within the grain interior distinctly reflects the significant imprint of the grain boundaries that disappeared during sintering.
@@
    Simulation results reveal two distinct types of GB cores. The first type, with nearly symmetric concentration profiles at the GB, exhibits stronger solute drag effects, resulting in a much lower moving velocity and a tendency to remain pinned in position. The second type of GB core shows asymmetric concentration profiles and a higher moving velocity, contributing predominantly to grain coarsening.
@@
    Abnormal grain growth resulting from solute drag effects leads to a large variation of grain boundary potentials. Typically, smaller grains exhibit higher grain boundary potentials at the later stages of grain growth and form more blocking GBs in term of ionic conductivity. When considering the total conductivity of an electrolyte material, the electrical current needs to find a ”detour” around these blocking GBs by passing through the surrounding larger grains, with less blocking GBs. It implies that to increase the conductivity one may also optimize the microstructure so that blocking GBs can be circumvented more easily. In contrast, if blocking GBs are desired, such as in capacitors, the grain size should be kept small to prevent the development of detour paths.
The proposed phase-field model offers naturally a series of promising extensions. We can e.g. incorporate additional species, such as neutral dopants, electrons, and electron holes, into the current phase-field model to quantitatively investigate the formation of the SCL. Furthermore, GB orientation plays a crucial role in SCL formation, and both the interface energy and the defect chemistry parameters $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\omega_{c}N_{\text{V}_{\text{O}}^{..},\text{c}}$ depend on GB orientation [49]. These complexities can be included in the phase-field model in the near future. Moreover, we assume in this paper only the $\text{Fe}^{3+}$ and ignore $\text{Fe}^{4+}$, which is reasonable for high temperature scenarios. But, the situation regarding solute drag effect changes at low temperature when a significant amount of $\text{Fe}^{4+}$ is present. The $\text{Fe}^{3+}$ ions in the GB core can not easily follow the movement of the GB core due to their low diffusivity. In contrast, the rapid movement of electron holes can adjust the local $\text{Fe}^{3+}$/$\text{Fe}^{4+}$ ratio and establish an equilibrium charge distribution. The solute drag effect may differ when a significant amount of $\text{Fe}^{4+}$ is considered.

### v3

@@ 6 Conclusion
    With the consideration of distinct segregation energies and available site densities of oxygen vacancy and acceptor dopant in bulk and GB core, the proposed phase-field model reproduces a series of benchmark results at equilibrium of Fe-doped STO bicrystal. Simulation results, including the symmetric SCLs, grain boundary potentials, concentrations of oxygen vacancies and acceptor dopants in the GB core and the bulk phase as well as the electrostatic distributions, show good agreement with the analytical predictions in the MS and GC models.
@@
    Simulations are carried out comprehensively for a GB core moving with constant velocities under quasi-equilibrium state. At low core velocity, the asymmetric SCL forms. With increasing GB core velocity, the dopant segregation diminishes. However, the movement of oxygen vacancy within the GB core only breaks at extremely high velocity due to the high diffusivity of oxygen vacancy. In addition, the electrostatic potential in the shrinking grain becomes higher than that in the growing grain.
@@
    Skewed grain size distribution is observed even without distinguishing any grain orientation or considering any anisotropic GB mobility in the phase-field simulations at 1623K. It indicates that the solute drag effect alone can lead to skewed grain size distribution that do not follow the log-normal distribution in the normal grain growth. As the formation energy differences of oxygen vacancies and acceptor dopants ($\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$) become more negative, the grain growth velocity decreases. This trend is also observed with increasing dopant concentration ($c_{\text{dop},\text{b}}$). Notably, the dopant concentration profile within the grain interior distinctly reflects the significant imprint of the grain boundaries that disappeared during sintering.
@@
    Simulation results reveal two distinct types of GB cores. The first type, with nearly symmetric concentration profiles at the GB, exhibits stronger solute drag effects, resulting in a much lower moving velocity and a tendency to remain pinned in position. The second type of GB core shows asymmetric concentration profiles and a higher moving velocity, contributing predominantly to grain coarsening. The coexistence of these grain boundaries results in grain growth behavior that deviates from the typical normal grain growth pattern..
@@
    To improve the sintering process, reducing the segregation energy of oxygen vacancies proves to be an effective approach. Lowering the dopant segregation energy and bulk concentration also mitigate solute drag effects. However, even at low dopant segregation energy and bulk dopant concentration, the skewed grain size distribution can still occur.
  • 6.
    The skewed grain size distribution resulting from solute drag effects leads to a large variation of grain boundary potentials. Typically, smaller grains exhibit higher grain boundary potentials at the later stages of grain growth and form more blocking GBs in term of ionic conductivity. When considering the total conductivity of an electrolyte material, the electrical current needs to find a ”detour” around these blocking GBs by passing through the surrounding larger grains, with less blocking GBs. It implies that to increase the conductivity one may also optimize the microstructure so that blocking GBs can be circumvented more easily. In contrast, if blocking GBs are desired, such as in capacitors, the grain size should be kept small to prevent the development of detour paths.
The proposed phase-field model offers naturally a series of promising extensions. First, we can e.g. incorporate additional species, such as the electrons, holes and the dopants with different valance states, into the current phase-field model to quantitatively investigate the formation of the SCL. Second, GB orientation plays a crucial role in SCL formation and different grain growth patterns, and both the interface energy and the defect chemistry parameters $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\omega_{c}N_{\text{V}_{\text{O}}^{..},\text{c}}$ depend on GB orientation de2019effect . Third, the assumptions of dilute and non-interacting defects become invalid at high defect concentrations, where defect–defect interactions must be accounted for. Various approaches have been proposed to address this complexity. One approach involves using a regular solution model to formulate the interaction energy contribution lund2021thermodynamically , vikrant2020electrochemical , vikrant2020electrochemicallyActa . Another approach, proposed by Mebane and De Souza (MDS) mebane2015generalised, provides expressions for the electrochemical potentials of oxygen vacancy and acceptor dopant while incorporating defect–defect interactions. However, MDS model does not include an explicit free energy formulation, which is essential for direct integration into phase-field models. Moreover, we assume in this paper only the $\text{Fe}^{3+}$ and ignore $\text{Fe}^{4+}$, which is reasonable for high temperature scenarios. But, the situation regarding solute drag effect changes at low temperature when a significant amount of $\text{Fe}^{4+}$ is present. The $\text{Fe}^{3+}$ ions in the GB core can not easily follow the movement of the GB core due to their low diffusivity. In contrast, the rapid movement of electron holes can adjust the local $\text{Fe}^{3+}$$\text{Fe}^{4+}$ ratio and establish an equilibrium charge distribution. The solute drag effect may differ when a significant amount of $\text{Fe}^{4+}$ is considered.